# Towards Global AI Inclusivity: A Large-Scale Multilingual Terminology Dataset (GIST)

### v3

@@ 5 Dataset Coverage Assessment
Eval Set Model Metric Arabic Chinese French Japanese Russian
60-60 aya-expanse-8B BLEU 20.11 + 1.23 + 0.18 27.31 + 1.33 + 0.24 33.05 + 2.46 + 0.20 14.59 + 0.61 + 0.32 16.59 + 1.59 - 0.05
COMET 81.96 + 0.71 - 0.52 83.43 + 1.57 + 0.08 81.83 + 1.06 - 0.11 88.54 + 0.32 - 0.01 82.27 + 0.69 - 2.02
aya-23-8B BLEU 19.98 + 0.54 - 0.21 26.08 + 0.47 + 0.39 33.85 + 2.28 - 0.11 15.06 + 0.87 + 0.36 15.77 + 1.05 + 0.37
COMET 84.02 + 0.81 - 0.24 85.12 + 0.58 + 0.38 82.40 + 0.94 - 0.15 87.92 + 0.50 + 0.09 81.91 + 0.40 - 2.26
gpt-4o-mini BLEU 23.58 + 1.07 - 0.00 32.64 + 1.60 + 0.66 40.80 + 3.08 + 0.50 21.46 + 0.64 + 0.19 17.25 + 1.07 - 0.13
COMET 85.77 + 0.69 - 0.44 87.30 + 0.48 + 0.26 84.56 + 0.68 - 0.04 89.96 + 0.14 + 0.01 83.68 + 0.38 - 2.29
nllb BLEU 22.38 + 1.37 + 0.64 17.29 + 1.92 + 1.02 34.93 + 2.86 + 0.21 6.19 + 2.42 + 0.53 17.30 + 1.54 + 1.07
COMET 83.52 + 0.83 - 0.45 78.22 + 2.95 + 0.73 82.83 + 1.00 - 0.19 77.82 + 3.80 + 0.39 81.41 + 0.97 - 1.54
seamless BLEU 23.13 + 1.16 - 0.03 26.26 + 0.97 + 0.80 40.04 + 2.08 - 0.57 14.56 + 0.74 + 0.05 17.18 + 1.71 + 1.17
COMET 84.07 + 0.94 - 0.38 83.44 + 1.48 + 0.50 83.86 + 0.78 - 0.07 85.05 + 1.06 + 0.16 82.33 + 0.56 - 1.87
AI Papers & Model Cards aya-expanse-8B BLEU 11.47 + 0.37 + 0.10 12.04 + 0.94 + 0.17 18.84 + 1.71 - 0.85 8.11 - 0.03 + 0.04 13.84 + 0.21 + 0.32
COMET 80.98 + 0.46 - 0.51 82.42 + 0.56 - 0.01 81.16 + 0.36 - 0.79 85.48 + 0.34 + 0.07 82.76 + 0.47 - 2.10
aya-23-8B BLEU 14.28 + 0.81 + 0.28 14.50 + 0.39 + 0.19 24.49 + 2.36 + 0.08 9.22 + 0.23 + 0.35 16.39 + 1.36 + 0.61
COMET 81.55 + 1.03 - 0.62 83.88 + 0.68 + 0.04 82.55 + 1.22 - 0.70 84.42 + 0.81 + 0.02 82.72 + 1.14 - 2.08
gpt-4o-mini BLEU 14.37 + 0.53 - 0.42 17.21 + 1.22 + 1.39 24.45 + 4.38 + 1.28 10.55 + 0.05 + 0.03 18.02 + 1.52 + 0.40
COMET 83.56 + 0.86 - 0.19 86.08 + 0.25 - 0.13 84.75 + 0.33 - 1.03 87.91 + 0.16 - 0.01 84.92 + 0.44 - 2.02
nllb BLEU 15.42 - 0.31 - 0.77 10.24 + 2.19 + 2.07 22.68 + 2.73 + 0.90 8.24 + 1.08 + 0.88 19.18 - 0.10 - 0.24
COMET 81.23 + 1.28 - 0.50 80.19 + 1.61 + 0.38 78.70 + 3.81 - 1.59 83.05 + 1.71 + 0.70 80.46 + 2.80 - 2.38
seamless BLEU 15.38 + 1.09 + 0.45 13.67 + 1.10 + 0.73 24.34 + 5.21 + 1.49 9.42 + 0.56 + 0.42 18.43 + 0.95 + 0.35
COMET 81.96 + 1.18 - 0.39 80.70 + 2.18 + 0.16 83.76 + 0.97 - 0.94 83.70 + 0.88 + 0.10 83.12 + 1.50 - 1.79
@@
Table 2: Experiment results on two evaluation sets across five models and five languages using BLEU and COMET. The first value represents the scores for direct translation using the models. The second value indicates the improvement or decline when applying the prompting-powered refinement method compared to direct translation. The third value indicates the improvement or decline when using the word alignment method compared to direct translation.
@@ 6.1 Methods
We apply the word alignment approach introduced by Dou and Neubig (2021) to identify and substitute term translations within the output. We use the multilingual BERT base model (Devlin et al., 2018) to tokenize both source and target sentences, and then processes these inputs through its hidden layers to produce contextualized embeddings. We determine alignments by computing dot-product similarities between source and target token embeddings. High-confidence alignments are then filtered using a threshold of 1e-4, and subword alignments are aggregated to generate word-level mappings. Lastly, we directly replace identified term translations with those from GIST.
@@ 6.3 Experiment Results
The word alignment method demonstrates mixed performance, showing improvements for Chinese and Japanese translations but leading to declines for Arabic, French, and Russian. This discrepancy is due to linguistic differences: Chinese and Japanese exhibit minimal morphological changes, allowing straightforward substitution of terminology without disrupting surrounding syntax; In contrast, languages like Arabic often require agreement in gender, number, and syntactic roles, making noun replacement more challenging and prone to errors. To validate this claim, we conduct a post-hoc prompting after word alignment using GPT-4o-mini to render the translations morphologically coherent and correct. Details of the prompting process are provided in Appendix B. As shown in Appendix C, this improves translation scores to align with the performance of the prompting-powered method across all languages and models.

### v4

@@ 5 Dataset Coverage Assessment
Model Metric Arabic Chinese French Japanese Russian
Evaluation Set: 60-60
aya-expanse BLEU 20.11 + 1.23 + 0.18 + 1.24 27.31 + 1.33 + 0.24 + 0.53 33.05 + 2.46 + 0.20 + 1.97 14.59 + 0.61 + 0.32 + 1.08 16.59 + 1.59 - 0.05 + 0.92
COMET 81.96 + 0.71 - 0.52 + 1.27 83.43 + 1.57 + 0.08 + 1.63 81.83 + 1.06 - 0.11 + 1.35 88.54 + 0.32 - 0.01 + 0.38 82.27 + 0.69 - 2.02 + 0.98
aya-23-8B BLEU 19.98 + 0.54 - 0.21 + 0.58 26.08 + 0.47 + 0.39 + 0.45 33.85 + 2.28 - 0.11 + 2.48 15.06 + 0.87 + 0.36 + 1.24 15.77 + 1.05 + 0.37 + 0.84
COMET 84.02 + 0.81 - 0.24 + 1.05 85.12 + 0.58 + 0.38 + 0.93 82.40 + 0.94 - 0.15 + 1.26 87.92 + 0.50 + 0.09 + 0.70 81.91 + 0.40 - 2.26 + 0.84
gpt-4o-mini BLEU 23.58 + 1.07 - 0.00 + 0.50 32.64 + 1.60 + 0.66 + 1.48 40.80 + 3.08 + 0.50 + 2.38 21.46 + 0.64 + 0.19 + 0.94 17.25 + 1.07 - 0.13 + 0.56
COMET 85.77 + 0.69 - 0.44 + 0.61 87.30 + 0.48 + 0.26 + 0.44 84.56 + 0.68 - 0.04 + 0.62 89.96 + 0.14 + 0.01 + 0.14 83.68 + 0.38 - 2.29 + 0.47
nllb BLEU 22.38 + 1.37 + 0.64 + 1.21 17.29 + 1.92 + 1.02 + 2.40 34.93 + 2.86 + 0.21 + 3.23 6.19 + 2.42 + 0.53 + 2.95 17.30 + 1.54 + 1.07 + 1.51
COMET 83.52 + 0.83 - 0.45 + 1.31 78.22 + 2.95 + 0.73 + 3.67 82.83 + 1.00 - 0.19 + 1.43 77.82 + 3.80 + 0.39 + 4.80 81.41 + 0.97 - 1.54 + 1.79
seamless BLEU 23.13 + 1.16 - 0.03 + 2.35 26.26 + 0.97 + 0.80 + 2.75 40.04 + 2.08 - 0.57 + 1.69 14.56 + 0.74 + 0.05 + 2.38 17.18 + 1.71 + 1.17 + 1.22
COMET 84.07 + 0.94 - 0.38 + 1.26 83.44 + 1.48 + 0.50 + 2.49 83.86 + 0.78 - 0.07 + 1.05 85.05 + 1.06 + 0.16 + 1.74 82.33 + 0.56 - 1.87 + 1.21
Evaluation Set: AI Papers & Model Cards
aya-expanse BLEU 11.47 + 0.37 + 0.10 + 0.49 12.04 + 0.94 + 0.17 + 0.85 18.84 + 1.71 - 0.85 + 2.19 8.11 - 0.03 + 0.04 + 0.31 13.84 + 0.21 + 0.32 + 0.68
COMET 80.98 + 0.46 - 0.51 + 0.63 82.42 + 0.56 - 0.01 + 0.75 81.16 + 0.36 - 0.79 + 0.44 85.48 + 0.34 + 0.07 + 0.54 82.76 + 0.47 - 2.10 + 0.75
aya-23-8B BLEU 14.28 + 0.81 + 0.28 + 0.75 14.50 + 0.39 + 0.19 + 0.71 24.49 + 2.36 + 0.08 + 2.37 9.22 + 0.23 + 0.35 + 0.45 16.39 + 1.36 + 0.61 + 1.35
COMET 81.55 + 1.03 - 0.62 + 1.23 83.88 + 0.68 + 0.04 + 0.80 82.55 + 1.22 - 0.70 + 1.23 84.42 + 0.81 + 0.02 + 1.26 82.72 + 1.14 - 2.08 + 1.29
gpt-4o-mini BLEU 14.37 + 0.53 - 0.42 + 0.61 17.21 + 1.22 + 1.39 + 1.34 24.45 + 4.38 + 1.28 + 3.40 10.55 + 0.05 + 0.03 + 0.01 18.02 + 1.52 + 0.40 + 1.60
COMET 83.56 + 0.86 - 0.19 + 0.93 86.08 + 0.25 - 0.13 + 0.07 84.75 + 0.33 - 1.03 + 0.13 87.91 + 0.16 - 0.01 + 0.19 84.92 + 0.44 - 2.02 + 0.49
nllb BLEU 15.42 - 0.31 - 0.77 + 0.29 10.24 + 2.19 + 2.07 + 3.10 22.68 + 2.73 + 0.90 + 3.40 8.24 + 1.08 + 0.88 + 1.21 19.18 - 0.10 - 0.24 + 1.44
COMET 81.23 + 1.28 - 0.50 + 1.99 80.19 + 1.61 + 0.38 + 2.84 78.70 + 3.81 - 1.59 + 1.62 83.05 + 1.71 + 0.70 + 2.55 80.46 + 2.80 - 2.38 + 1.63
seamless BLEU 15.38 + 1.09 + 0.45 + 1.66 13.67 + 1.10 + 0.73 + 1.99 24.34 + 5.21 + 1.49 + 5.35 9.42 + 0.56 + 0.42 + 0.91 18.43 + 0.95 + 0.35 + 1.70
COMET 81.96 + 1.18 - 0.39 + 1.72 80.70 + 2.18 + 0.16 + 2.95 83.76 + 0.97 - 0.94 + 0.88 83.70 + 0.88 + 0.10 + 1.81 83.12 + 1.50 - 1.79 + 1.70
@@
Table 2: Evaluation results across five models and five languages using BLEU and COMET metrics. The first black value in each column represents direct translation scores. The second, third, and fourth values (colored red or green) indicate the relative change in performance when applying the prompting-powered refinement method, the word alignment method, and post-hoc prompting after word alignment for improved morphological and grammatical correctness, respectively, compared to direct translation.
@@ 6.1 Methods
We apply the word alignment approach introduced by Dou and Neubig (2021) to identify and substitute term translations within the output. We use the multilingual BERT base model (Devlin et al., 2018) to tokenize both source and target sentences, and then processes these inputs through its hidden layers to produce contextualized embeddings. We determine alignments by computing dot-product similarities between source and target token embeddings. High-confidence alignments are then filtered using a threshold of 1e-4, and subword alignments are aggregated to generate word-level mappings. Lastly, we directly replace identified term translations with those from GIST. We also conduct a post-hoc prompting after word alignment using GPT-4o-mini to render the translations morphologically coherent and correct. Details of the prompting process are provided in Appendix B.
@@ 6.3 Experiment Results
The word alignment method demonstrates mixed performance, showing improvements for Chinese and Japanese translations but leading to declines for Arabic, French, and Russian. This discrepancy is due to linguistic differences: Chinese exhibits minimal morphological changes, allowing straightforward substitution of terminology with limited disruption to surrounding syntax; In contrast, languages like Arabic often require agreement in gender, number, and syntactic roles, making noun replacement more complex and error-prone. By comparing the fourth and second values in each column of Table 2, we observe that the post-hoc prompting ablation enhances translation scores, aligning them with the performance of the prompting-powered method across all languages and models, thereby supporting our claims.